if E=1/2mv2 .where m=52.4+ or _0.2,v=25.6+ or_0.1 what is the percentage error in F

1
Expert's answer
2021-07-22T10:08:26-0400

"E=\frac{1}{2}mv^2"

Error

"\frac{∆E}{E}=\frac{∆m}{m}+2\frac{∆v}{v}"

"\frac{∆E}{E}=\frac{0.2}{52.4}+2\times \frac{0.1}{25.6}"

"\frac{∆E}{E}" =0.00381+0.00781=0.01162

"\frac{∆E}{E}=0.01162"

Percentage error


"\frac{∆E}{E}\times100=0.01162\times100=1.162%"

"\frac{∆E}{E}"%=1.162%